what is mean by electric constant?
electric constant=1/(magnetic constant * speed of light)

please explain this equation.

Grade:12

2 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

The electric constant0) is a constant.

it appears Coulomb's law can be given as:

 F = \frac{1}{4 \pi \varepsilon_0} \ \frac{|q| |Q|}{r^2}

\varepsilon_0 = \frac{1}{\mu_0 c^2}

here μ0 = 4π × 10−7 N/A2

\varepsilon_0 = \frac{10^7}{4\pi\,c^2} =  8.854\;187\;817... 10^{-12}
ε0=The permittivity of empty space
   = equal to 1 in centimeter-gram-second electrostatic units and to 107/4πc2 farads permeter         = 8.854 × 10-12 farad per meter 


Now you can win exciting gifts by answering the questions on
Discussion Forum. So help discuss any query on askiitians forum and
become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the
Askiitians Toolbar : Click here to download the toolbar..


harini desiraju
13 years ago

the terms electric constant and magnetic constant are introduced while defining coloumb's inverse square law.These are used as proportionality constants in the fields of electricity and magnetism respectively.I am not aware of the derivation,but can tell u the relation between 'c' and these constants. light is an electromagnetic wave.So,its speed should describe the nature of both electric and magnetic waves. Thus we can accept this relation.(c2=1/electric const. * magnetic const.)